OpenStudy (anonymous):

If the wheel is spun and each section is equally likely to stop under the pointer, determine the probability that the pointer lands on a number greater than 6, given that the color is red.

OpenStudy (anonymous):

the wheel

OpenStudy (anonymous):

wow :)

OpenStudy (anonymous):

the wheel

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

the wheel

OpenStudy (anonymous):

\[\frac{1}{3}\]

OpenStudy (anonymous):

i think its 1/12

OpenStudy (anonymous):

nice wheel.

OpenStudy (anonymous):

the wheel

OpenStudy (anonymous):

lol i posted a file with it..

OpenStudy (anonymous):

the wheel

OpenStudy (anonymous):

someone told me it was 1/12 thats why i posted it again...

OpenStudy (anonymous):

I think 1/3

OpenStudy (anonymous):

the wheel

OpenStudy (anonymous):

thats what i thought cuz there is 3 reds and only 1 thats over 6

OpenStudy (anonymous):

you are told that the color that came up was red. Since there are 3 red sections, and only one of them has a number greater than 6 (it's 7) - the probability is 1/3

OpenStudy (anonymous):

the wheel

OpenStudy (anonymous):

ok ty :)

OpenStudy (zarkon):

\[\frac{1/12}{3/12}=\frac{1}{3}\]

OpenStudy (amistre64):

P(n>6, given red) = P(n>6 and red)/P(red)

OpenStudy (amistre64):

1/12 is a number > 6 and red 3/12 are red 1/12 ---- = 1/3 3/12

OpenStudy (anonymous):

the wheel

OpenStudy (anonymous):

i think its 1/3 for answer